# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ 5 Quantized Nets Sample Complexity
Recall that the generalization bound in Theorems 3.2 depends on the effective sample complexity $\tilde{C}\triangleq-\log\left(\tilde{p}\right)$. In this section, we derive an upper bound on $\tilde{C}$ for quantized multi-layer Fully Connected (FC) Neural Networks (NNs) with a single binary output, with and without additional per-node scaling.
@@
For a depth $L$, widths $D=\left(d_{1},\dots,d_{L}\right)$, and activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a fully connected NN is a mapping $\bm{\theta}\mapsto h^{FC}_{\bm{\theta}}$ from parameters
$\left\{\bm{\theta}=\left\{\mathbf{W}^{\left(l\right)},\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L}\,\middle|\,\mathbf{W}^{\left(l\right)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{d_{l}}\right\}$
@@
$h_{\bm{\theta}}^{FC}\left(\mathbf{x}\right)$ $=\operatorname{sign}\!\left(\mathbf{W}^{\left(L\right)}f^{(L-1)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(L\right)}\right)\!.$
@@
As we will show, considering NNs in which each neuron is multiplied by a scaling parameter can significantly improve the bound. This architecture modification is common in empirical practices, e.g. batch-normalization (Ioffe and Szegedy, 2015), weight-normalization (Salimans and Kingma, 2016), and certain initializations (Zhang et al., 2019). Formally, we define a Scaled-neuron Fully Connected (SFC) NN as
@@
For a depth $L$, widths $D=\left(d_{1},\dots,d_{L}\right)$, and activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a scaled neuron fully connected neural network is a mapping $\bm{\theta}\mapsto h^{FC}_{\bm{\theta}}$ from parameters
$\bm{\theta}=\left\{\mathbf{W}^{\left(l\right)},\mathbf{b}^{\left(l\right)},\bm{\gamma}^{\left(l\right)}\right\}_{l=1}^{L}\,,$
where $\mathbf{W}^{\left(l\right)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{d_{l}},\bm{\gamma}^{\left(l\right)}\in\mathbb{R}^{d_{l}}\,,$ defined recursively, starting with $f^{(0)}\left(\mathbf{x}\right)=\mathbf{x}\,,$ then
$\forall l\in\left[L-1\right]\;f^{(l)}\left(\mathbf{x}\right)$ $=\sigma\left(\bm{\gamma}^{\left(l\right)}\odot{\mathbf{W}^{\left(l\right)}f^{(l-1)}\left(\mathbf{x}\right)}+\mathbf{b}^{\left(l\right)}\right)$
$h_{\bm{\theta}}^{FC}\left(\mathbf{x}\right)$ $=\operatorname{sign}\left(\mathbf{W}^{\left(L\right)}f^{(L-1)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(L\right)}\right).$
@@
We consider $Q$-quantized networks where each of the parameters is chosen from a fixed set $\mathcal{Q}\subset\mathbb{R}$ such that $0\in\mathcal{Q}$ and ${\left\lvert{\mathcal{Q}}\right\rvert}\leq Q$. This can be the set of integers $\{-\frac{Q}{2},-\frac{Q}{2}+1,\ldots,(\frac{Q}{2}-1)\}$ for even $Q$, or the set of numbers representable as $\log_{2}Q$-bit floats (for, e.g. $\log_{2}Q=32$). Fully connected quantized NNs thus have parameters $\bm{\theta}\in\mathcal{Q}^{M}$ corresponding to a complexity $C=M\log Q$ (from the classic log cardinality bound, see Appendix A.1).
We consider a teacher $h^{\star}=h_{\bm{\theta}^{\star}}$ that is a $Q$-quantized network of some depth $L$ and small widths $D^{\star}=(d^{\star}_{1},\ldots,d^{\star}_{L})$, and a wider student of the same depth $L$ but widths $D\gg D^{\star}$. For the student, we consider a uniform prior over $Q$-quantized parametrizations, i.e. $\bm{\theta}\sim\textrm{Uniform}\left(Q^{M(D)}\right)$. In other words, to generate $h_{\bm{\theta}}\sim\mathcal{P}$, each weight (and bias) in the NN is chosen independently and uniformly from $\mathcal{Q}$.
@@ 5.1 Main Results
For any activation function such that $\sigma(0)=0$, depth $L$, $Q$-quantized teacher with widths $D^{\star}$, student with widths $D>D^{\star}$, $d_{0}^{\star}\triangleq d_{0}$, and prior $\mathcal{P}$ uniform over $Q$-quantized parametrizations, we have that:
@@
And, by Theorem 3.2, $N=(\tilde{C}+3\log 2/\delta)/\epsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\epsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
Proof idea. The idea is simple and centers on counting the necessary number of constraints on the parameters of the student network to ensure it is TE. A fundamental illustration of this concept is provided in the caption of Figure 1. Full proof is in Appendix B.
@@ 5.2 Discussion: Comparing Sample Complexities
$C^{\star}=\left(\sum_{l=1}^{L}\left(d_{l}^{\star}d_{l-1}^{\star}+kd_{l}^{\star}\right)\right)\log Q\quad;\quad C=\left(\sum_{l=1}^{L}\left(d_{l}d_{l-1}+kd_{l}\right)\right)\log Q$
@@
Either way, the dominant term is the quadratic term $\sum_{l}d_{l}d_{l-1}$. Lacking other considerations, the sample complexity of learning with the student network would be $C$. However, we see here that thanks to the parametrization and prior, the student network implicitly adapts to the complexity of the teacher, with $\hat{C}\ll C$ when $D^{\star}\ll D$ and in any case $C^{\star}\leq\hat{C}^{\mathrm{}}\leq C$. With Vanilla FC Networks, the sample complexity $\hat{C}^{\mathrm{FC}}$, although smaller than $C$, is still significantly larger than the complexity $C^{\star}$ of the teacher, which is what we could have hoped for. The quadratic (dominant) terms in $\hat{C}^{\mathrm{FC}}$ are roughly geometric averages of terms from $C$ and $C^{\star}$, and so we have that, very roughly, $\hat{C}^{\mathrm{FC}}\approx\sqrt{CC^{\star}}$. We can improve this using scaling, which creates more redundancy since zero scales can deactivate entire units. Indeed, for Scaled Fully Connected Networks, we have that $\hat{C}^{\mathrm{SFC}}=C^{\star}+\sum_{l=1}^{L}d_{l}\ll\hat{C}^{\mathrm{FC}}\ll C$ (when $D^{\star}\ll D$). We still pay a bit for the width of the student, but only linearly instead of quadratically. In particular, even if the width of the student is quadratic in the width of the teacher, the sample complexity of learning by sampling random NNs is almost the same as that of using a much narrower teacher.
Minimum widths. We note that, just as in the description length example in the previous section, an explicit narrowness prior could have of course been fully adaptive to the width of the teacher and ensured learning with the ideal sample complexity $C^{\star}$. E.g., this is achieved by allowing the student to choose the width of each layer, and using the Occam rule
$\min_{D^{\prime}\leq D,\bm{\theta}\in\mathcal{Q}^{M(D^{\prime})}\text{with widths}D^{\prime}}M(D^{\prime})\;\;\text{s.t.~{}}\mathcal{L}_{S}(h_{\bm{\theta}})=0.$ (7)
@@
$\min_{\bm{\theta}\in\mathcal{Q}^{M(D)}}\left\|\bm{\theta}\right\|_{0}\;\;\text{s.t.~{}}\mathcal{L}_{S}(h_{\bm{\theta}})=0.$ (8)
The sparsity-inducing rule (8) would have the following bound for the effective sample complexity.
$\hat{C}^{\textrm{sparse}}\triangleq O\left(M\left(D^{\star}\right)\log\left(M\left(D^{\star}\right)+d_{0}\right)+M\left(D^{\star}\right)\log Q\right)\,.$ (9)
See Appendix D for a derivation of this equation. Note that $\hat{C}^{\textrm{sparse}}$ in (9) does not depend on the size of the student, but rather only on the size of the teacher, which is smaller, that is $M\left(D^{\star}\right)\ll M\left(D\right)$. For comparison, recall our previous bound for posterior sampling, in the case of FC scaled-neuron networks from (6):
$\hat{C}^{\mathrm{SFC}}=\left(\sum_{l=1}^{L}\left(d_{l}^{\star}d_{l-1}^{\star}+2d_{l}\right)\right)\log Q=O\left(C^{\star}+\sum_{l=1}^{L}2d_{l}\log Q\right)\,.$
@@
Minimum norm. One might also ask whether a similar adaptation to teacher width can be obtained by regularizing the norm of the weights. Indeed, Neyshabur et al. (2015); Golowich et al. (2018) obtain sample complexity bounds that depend only on the $\ell_{2}$ norm $\left\|\bm{\theta}\right\|$ of the learned network, without any dependence on the width of the student (but with an exponential dependence on depth!). These guarantees are not directly applicable in our setting, since applying them to get guarantees on the misclassification error we study here requires bounding the margin. Even with a discrete teacher, without further assumptions on the input distribution, we cannot ensure a margin. If we did consider only integer inputs and integer weights, we could at least ensure a margin of 1. In this case: on one hand, the norm-based sample complexity would scale as $Q^{O(L)}$, i.e. exponential in the $L\log Q$ dependence of $\hat{C}^{\mathrm{}}$. On the other hand, the norm-based guarantee would not depend at all on the student widths $D$, while even $\hat{C}^{\mathrm{SFC}}$ increases linearly with the student widths.
@@ 5.3 Extension to convolutional neural networks
In Appendix B, we extend Theorem 5.3 to convolutional neural networks (CNN) and convolutional neural networks where each neuron input is multiplied by a learned parameter (SCNN). Specifically, we show that for quantized convolutional networks, we get similar bounds on the sample complexity $\hat{C}$ with channel numbers substituting layer widths.
@@
where $d_{s}$ is the number of neurons in the last convolutional layer (i.e., the width of the last layer which is a fully connected layer with a single output), $k_{l}$ and $c_{l}$ are the $l^{\text{th}}$ layer kernel size and number of channels. See Appendix B for precise definitions of the model and statement of the result. Importantly, similar to the fully connected architecture we observe that $\hat{C}^{\mathrm{SCNN}}$ is again near to the teacher complexity, with only a weak dependency on the student channel numbers. The proof here is analogous to the proof of Theorem 5.3, where neurons are replaced with channels.

### v2

@@ 4 Quantized Nets Sample Complexity
Recall that the generalization bound in Lemma 3.2 depends on the effective sample complexity $\tilde{C}\triangleq-\log\left(\tilde{p}\right)$. In this section, we derive an upper bound on $\tilde{C}$ for quantized multi-layer Fully Connected (FC) Neural Networks (NNs) with a single binary output, with and without additional per-node scaling.
@@
For a depth $L$, widths $D=\left(d_{1},\dots,d_{L}\right)$, and activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a fully connected NN is a mapping $\boldsymbol{\theta}\mapsto h^{FC}_{\boldsymbol{\theta}}$ from parameters
$\left\{\boldsymbol{\theta}=\left\{\mathbf{W}^{\left(l\right)},\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L}\,\middle|\,\mathbf{W}^{\left(l\right)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{d_{l}}\right\}$
defined recursively, starting with $f^{(0)}\left(\mathbf{x}\right)=\mathbf{x}\,,$ as
@@
$h_{\boldsymbol{\theta}}^{\mathrm{FC}}\left(\mathbf{x}\right)$ $=\operatorname{sign}\!\left(\mathbf{W}^{\left(L\right)}f^{(L-1)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(L\right)}\right)\!.$
@@
As we will show, considering NNs in which each neuron is multiplied by a scaling parameter can significantly improve the bound. This architecture modification is common in empirical practices, e.g. batch-normalization (Ioffe & Szegedy, 2015), weight-normalization (Salimans & Kingma, 2016), and certain initializations (Zhang et al., 2019). We formally define this model in the following definition.
@@
For a depth $L$, widths $D=\left(d_{1},\dots,d_{L}\right)$, and activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a scaled neuron fully connected neural network is a mapping $\boldsymbol{\theta}\mapsto h^{\mathrm{SFC}}_{\boldsymbol{\theta}}$ from parameters
$\boldsymbol{\theta}=\left\{\mathbf{W}^{\left(l\right)},\mathbf{b}^{\left(l\right)},\boldsymbol{\gamma}^{\left(l\right)}\right\}_{l=1}^{L}\,,$
where $\mathbf{W}^{\left(l\right)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{d_{l}},\boldsymbol{\gamma}^{\left(l\right)}\in\mathbb{R}^{d_{l}}\,,$ defined recursively, starting with $f^{(0)}\left(\mathbf{x}\right)=\mathbf{x}\,,$ as
$\forall l\in\left[L-1\right]\;f^{(l)}\left(\mathbf{x}\right)$ $=\sigma\left(\boldsymbol{\gamma}^{\left(l\right)}\odot{\mathbf{W}^{\left(l\right)}f^{(l-1)}\left(\mathbf{x}\right)}+\mathbf{b}^{\left(l\right)}\right)$
$h_{\boldsymbol{\theta}}^{\mathrm{SFC}}\left(\mathbf{x}\right)$ $=\operatorname{sign}\left(\mathbf{W}^{\left(L\right)}f^{(L-1)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(L\right)}\right).$
@@
We consider $Q$-quantized networks where each of the parameters is chosen from a fixed set $\mathcal{Q}\subset\mathbb{R}$ such that $0\in\mathcal{Q}$ and ${\left\lvert{\mathcal{Q}}\right\rvert}\leq Q$. This can be the set of integers $\{-\frac{Q}{2},-\frac{Q}{2}+1,\ldots,(\frac{Q}{2}-1)\}$ for even $Q$, or the set of numbers representable as $\log_{2}Q$-bit floats (for, e.g. $\log_{2}Q=32$). Fully connected quantized NNs thus have parameters $\boldsymbol{\theta}\in\mathcal{Q}^{M}$ corresponding to a complexity $C=M\log Q$ (from the classic log cardinality bound, see Appendix B.1).
We consider a teacher $h^{\star}=h_{\boldsymbol{\theta}^{\star}}$ that is a $Q$-quantized network of some depth $L$ and small widths $D^{\star}=(d^{\star}_{1},\ldots,d^{\star}_{L})$, and a wider student of the same depth $L$ but widths $D>D^{\star}$. For the student, we consider a uniform prior over $Q$-quantized parameterizations, i.e. $\boldsymbol{\theta}\sim\textrm{Uniform}\left(Q^{M(D)}\right)$. In other words, to generate $h_{\boldsymbol{\theta}}\sim\mathcal{P}$, each weight (and bias) in the NN is chosen independently and uniformly from $\mathcal{Q}$.
@@ 4.1 Main Results
For any activation function such that $\sigma(0)=0$, depth $L$, $Q$-quantized teacher with widths $D^{\star}$, student with widths $D>D^{\star}$, $d_{0}^{\star}\triangleq d_{0}$, and prior $\mathcal{P}$ uniform over $Q$-quantized parameterizations, we have that:
@@
And, by Lemma 3.2, $N=(\tilde{C}+3\log 2/\delta)/\varepsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\varepsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
Remark 4.4.
From Corollary 3.3 we deduce that for a given training set, the volume of ‘bad’ ($\varepsilon$) interpolators is ‘small’ ($\delta$) with high probability ($1-\delta$) with this sample complexity.
Proof idea. The idea is simple and centers on counting a sufficient number of constraints on the parameters of the student network to ensure it is TE. A fundamental illustration of this concept is provided in the caption of Figure 1. Full proof is in Appendix C.
@@ 4.2 Discussion: Comparing Sample Complexities
$C^{\star}=\left(\sum_{l=1}^{L}\left(d_{l}^{\star}d_{l-1}^{\star}+kd_{l}^{\star}\right)\right)\log Q$
$C=\left(\sum_{l=1}^{L}\left(d_{l}d_{l-1}+kd_{l}\right)\right)\log Q$
@@
Either way, the dominant term is the quadratic term $\sum_{l}d_{l}d_{l-1}$. Lacking other considerations, the sample complexity of learning with the student network would be $C$. However, we see here that thanks to the parameterization and prior, the student network implicitly adapts to the complexity of the teacher, with $\hat{C}\ll C$ when $D^{\star}\ll D$ and in any case $C^{\star}\leq\hat{C}^{\mathrm{}}\leq C$. With Vanilla FC Networks, the sample complexity $\hat{C}^{\mathrm{FC}}$, although smaller than $C$, is still significantly larger than the complexity $C^{\star}$ of the teacher, which is what we could have hoped for. The quadratic (dominant) terms in $\hat{C}^{\mathrm{FC}}$ are roughly geometric averages of terms from $C$ and $C^{\star}$, and so we have that, very roughly, $\hat{C}^{\mathrm{FC}}\approx\sqrt{CC^{\star}}$. We can improve this using scaling, which creates more redundancy since zero scales can deactivate entire units. Indeed, for Scaled Fully Connected Networks, we have that $\hat{C}^{\mathrm{SFC}}=C^{\star}+\sum_{l=1}^{L}d_{l}\ll\hat{C}^{\mathrm{FC}}\ll C$ (when $D^{\star}\ll D$). We still pay a bit for the width of the student, but only linearly instead of quadratically. In particular, even if the width of the student is quadratic in the width of the teacher, the sample complexity of learning by sampling random NNs is almost the same as that of using a much narrower teacher.
Minimum widths. We note that, just as in the minimum description length example in the previous section, an explicit narrowness prior could have of course been fully adaptive to the width of the teacher and ensured learning with the ideal sample complexity $C^{\star}$. E.g., this is achieved by allowing the student to choose the width of each layer, and using the Occam rule
$\min_{D^{\prime}\leq D,\boldsymbol{\theta}\in\mathcal{Q}^{M(D^{\prime})}\text{with widths}\,D^{\prime}}M(D^{\prime})\;\;\text{s.t.~{}}\mathcal{L}_{S}(h_{\boldsymbol{\theta}})=0.$ (8)
@@
$\min_{\boldsymbol{\theta}\in\mathcal{Q}^{M(D)}}\left\|\boldsymbol{\theta}\right\|_{0}\;\;\text{s.t.~{}}\mathcal{L}_{S}(h_{\boldsymbol{\theta}})=0.$ (9)
The sparsity-inducing rule (9) would have the following bound for the effective sample complexity
$\hat{C}^{\textrm{sparse}}\triangleq O\left(M\left(D^{\star}\right)\left(\log\left(M\left(D^{\star}\right)+d_{0}\right)+\log Q\right)\right).$ (10)
See Appendix E for a derivation of this equation. Note that $\hat{C}^{\textrm{sparse}}$ in (10) does not depend on the size of the student, but rather only on the size of the teacher, which is smaller, that is $M\left(D^{\star}\right)\ll M\left(D\right)$. For comparison, recall our previous bound for posterior sampling, in the case of FC scaled-neuron networks from (7):
$\hat{C}^{\mathrm{SFC}}$ $=\left(\sum_{l=1}^{L}\left(d_{l}^{\star}d_{l-1}^{\star}+2d_{l}\right)\right)\log Q$
$=O\left(C^{\star}+\sum_{l=1}^{L}2d_{l}\log Q\right)\,.$
@@
Minimum norm. One might also ask whether a similar adaptation to teacher width can be obtained by regularizing the norm of the weights. Indeed, (Neyshabur et al., 2015; Golowich et al., 2018) obtained sample complexity bounds that depend only on the $\ell_{2}$ norm $\left\|\boldsymbol{\theta}\right\|$ of the learned network, without any dependence on the width of the student (but with an exponential dependence on depth!). These guarantees are not directly applicable in our setting, since applying them to get guarantees on the misclassification error we study here requires bounding the margin. Even with a discrete teacher, without further assumptions on the input distribution, we cannot ensure a margin. If we did consider only integer inputs and integer weights, we could at least ensure a margin of 1. In this case: on one hand, the norm-based sample complexity would scale as $Q^{O(L)}$, i.e. exponential in the $L\log Q$ dependence of $\hat{C}^{\mathrm{}}$. On the other hand, the norm-based guarantee would not depend at all on the student widths $D$, while even $\hat{C}^{\mathrm{SFC}}$ increases linearly with the student widths.
@@ 4.3 Extension to convolutional neural networks
In Appendix C, we extend Theorem 4.3 to convolutional neural networks (CNN) and convolutional neural networks where each channel is multiplied by a learned parameter (SCNN). Specifically, we show that for quantized convolutional networks, we get similar bounds on the sample complexity $\hat{C}$ with channel numbers substituting layer widths.
@@
where $d_{s}$ is the number of neurons in the last convolutional layer (i.e., the width of the last layer which is a fully connected layer with a single output), $k_{l}$ and $c_{l}$ are the $l^{\text{th}}$ layer’s kernel size and number of channels. See Appendix C for precise definitions of the model and statement of the result. Importantly, similar to the fully connected architecture we observe that $\hat{C}^{\mathrm{SCNN}}$ is again close to the teacher’s complexity, with only a weak dependence on the student’s channel numbers. The proof here is analogous to the proof of Theorem 4.3, where neurons are replaced with channels.
Implications to realistic benchmarks. Currently, no NN has achieved zero test error on real-world datasets, even for simple ones like MNIST. Therefore, the size of the teacher NN is unknown to us when facing practical applications. However, we can use the dimensions of common NN architectures and datasets to approximate the necessary size of the teacher to obtain meaningful generalization bounds using our results. For example, suppose that we substitute the actual sizes of a training set and network into our bound
$N=\frac{\hat{C}+3\log{\left(\frac{2}{\delta}\right)}}{\varepsilon}.$
We can deduce the size of the teacher required to satisfy this equation. For example, using the size of the ImageNet dataset, $\delta=0.05$, and the actual test error of some CNNs, we can estimate the required size of the teacher. For simplicity of calculation, we assume that each layer of the teacher NN has exactly $\alpha$ channels compared to the same layer in the student NN, where $10<\alpha<$ 1. In the following table, we show the required width reduction $\alpha$ and the number of parameters in the resulting teacher NN. We use the smallest quantization level (2bit) for which the NN accuracy remains near the FP32 accuracy (less than  0.5% degradation, from (Liu et al., 2022b)).
Architecture $\varepsilon$ $\alpha$parameters
ResNet18 0.3 0.125 $\sim 241k$
ResNet50 0.25 0.05 $\sim 159k$
Table 1: Approximate relative width reduction and number of parameters in the teacher required to obtain meaningful bounds using standard ResNet architectures and the ImageNet dataset.
Table 1 shows that our bound is consistent with a narrow teacher of non-trivial size and widths. For example, in ResNet18 the resulting channel numbers in the teacher layers are [3,8,8,16,16,32.32,64,.64]388161632.3264.64[3,8…,8,16,...16,32….32,64,....64][ 3 , 8 … , 8 , 16 , … 16 , 32 … .32 , 64 , … .64 ], where ‘3’ counts the input channels and the rest are the following hidden layers. This architecture can implement highly complex non-linear functions. Note that we cannot directly validate the existence of this teacher, as standard optimization methods may not be capable of finding such a solution efficiently, even if it exists (without over-parameterization, it is much harder to find global minima).